OpenStudy (abbycross167):

Can someone please help me with this question? What would the y-intercept of the equation y = -(3/4)x be? I would really like someone to show me how to do this... please

OpenStudy (anonymous):

to calculate for the y-intercept of the function, you need to substitute x=0 and plug it into the equation

OpenStudy (anonymous):

what will you get if x is 0?

OpenStudy (abbycross167):

-(3/4)*0= 0? right? @Data_LG2

OpenStudy (anonymous):

Yes, the function will pass through the origin.. so the y-int or x-int is (0,0) :)

OpenStudy (anonymous):

Things to remember: If you are looking for the y-int, always sub x=0 into the equation to solve for y. The y-int will be (0,y). If you are looking for the x-int, always sub y=0 into the equation to solve for x. The x-int will be (x,0) Hope this helps you in the future ^_^

OpenStudy (abbycross167):

@Data_LG2 Can I show you my answer to this question? My teacher said I got it incorrect

OpenStudy (anonymous):

sure

OpenStudy (abbycross167):

ok my answer~ The intercept is 0, y=3/4 of 0=0. The line goes through the point (0,0)

OpenStudy (anonymous):

Hmm.. Maybe you can just fix the phrasing. y is not 3/4 when x=0 but you wrote y=3/4 so probably your teacher assumed that your y-intercept is 3/4 not 0

OpenStudy (anonymous):

Try saying like this: The y-intercept is at origin (0,0) since graph will pass the y-axis at this point.

OpenStudy (abbycross167):

ok I will thank you soo much for helping!!

OpenStudy (anonymous):

no problem, see you around :)